OpenStudy (anonymous):

please help! 3n+6/3=15 N= ?

OpenStudy (anonymous):

i did my work and came up with 13/3 which goes into 4.3 repeating.. but i dont know how to write in the answer box or if its correct

OpenStudy (anonymous):

will medal and fan!

OpenStudy (anonymous):

sorry dont hnow it D:

OpenStudy (anonymous):

\[\frac{ 3n+6 }{ 3 }\] =15 is the question i got \[\frac{ 13 }{ 3}\]

OpenStudy (anonymous):

nevermind i inserted 4.3 and i was correct. thanks for trying